solve for x: log base 3, x squared, equals two log base 3, 4, minus 4 log basdse 3, 5
Answer by edjones(7569)

log[3](x^2)=2log[3](4)-4log[3](5)
log[3](x^2)=log[3](16)-log[3](625)
3^(log[3](x^2))=3^(log[3](16))/3^(log[3](625))
x^2=16/625
x=4/25, x=-4/25
.
Ed